OpenStudy (anonymous):

factors of 9x^2+3x+1?

OpenStudy (anonymous):

You have no factors.

OpenStudy (anonymous):

\[9x^2+3x+1=\Big(3x+\frac{1}{2}+\frac{\sqrt{3}}{2}i\Big)\Big(3x+\frac{1}{2}-\frac{\sqrt{3}}{2}i\Big)\]

OpenStudy (anonymous):

there is no factor..if you want to find the rrots or the value of x..you must go with the process of completing the square or use the quadratic formula..